### v1

@@ 8. Euler characteristics of perverse sheaves on ${\mathscr{M}}_{g,n}$
There is no loss in assuming that $g_{1}\geq g_{2}\geq\ldots\geq 0$. Let $<_{L}$ denote the lexicographic order on the set of the vectors $(g_{1},g_{2},\ldots)$. For example,
$(1,1,1)<_{L}(2,1)<_{L}(3)$
This is a well ordering, so we can use it to do induction. Fix $N\geq 3$, and let $X={}_{N}{\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times{}_{N}{\mathscr{M}}_{g_{r},n_{r}}^{c}$. It is enough to prove that
(8.4.1) $\chi(X,K)\geq 0$
for $K\in\operatorname{Perv}(X)$. Suppose that $g_{1}=1$, i.e. all $g_{i}=1$. Then we have an affine open immersion $j:X\to\prod{}_{N}{\mathscr{A}}_{1,n_{i}}$. Therefore $j_{*}K$ is perverse. Consequently theorem 5.2 implies
$\chi(X,K)=\chi(\prod{}_{N}{\mathscr{A}}_{1,n_{i}},j_{*}K)\geq 0$
Therefore we may assume that $g_{1}>1$. We prove (8.4.1) by induction on $(g_{1},g_{2},\ldots)$. For each $g_{i}\geq 2$ and decompositions $g_{i}=g_{1}^{\prime}+g_{2}^{\prime}$, with $g_{i}^{\prime}>0$ and $n_{i}=n_{1}^{\prime}+n_{2}^{\prime}$,, we have a map
$id\times\ldots gl\ldots:{}_{N}{\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times({}_{N}{\mathscr{M}}_{g_{1}^{\prime},n_{1}^{\prime}}^{c}\times{}_{N}{\mathscr{M}}_{g_{2}^{\prime},n_{2}^{\prime}}^{c})\times\ldots\to X$
Let $Z=X-\prod{}_{N}{\mathscr{M}}_{g_{i},n_{i}}$. When $g_{1}>1$, $Z$ is a union of images of the above gluing maps.
We claim that $\chi(Z,K^{\prime\prime})\geq 0$ for $K^{\prime\prime}\in\operatorname{Perv}(Z)$. Since $\operatorname{Perv}(Z)$ is Artinian, we can assume that $K^{\prime\prime}$ is simple. Then $K^{\prime\prime}$ is the image of a perverse sheaf $L$ in some $\prod{}_{N}{\mathscr{M}}_{g_{i}^{\prime},n_{i}^{\prime}}$, with $(g_{1}^{\prime},g_{2}^{\prime},\ldots)<_{L}(g_{1},g_{2},\ldots)$. Therefore the claim follows by induction.
By taking products of $\tau$, we get a morphism that we also denote by the same letter
$\tau:X\to\prod{}_{N}{\mathscr{A}}_{g_{i},n_{i}}$
Let $Y=\tau(X)$. Theorem 5.2 shows that $\chi(Y,K^{\prime})\geq 0$ for any $K^{\prime}\in\operatorname{Perv}(Y)$. This fact together with the claim and theorem 6.1 implies (8.4.1).
@@
Let ${\mathscr{M}}^{c}={\mathscr{M}}_{g_{1},n_{1}}^{c}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}^{c}$. The embedding $j:M\to M^{c}$ is affine, since the complement is a divisor. Therefore $j_{*}P\in\operatorname{Perv}({\mathscr{M}}^{c})$. Consequently,
$\chi^{orb}({\mathscr{M}},P)=\chi^{orb}({\mathscr{M}}^{c},j_{*}P)\geq 0$
We note that the previous theorem is false when the $g_{i}=0$. For example, ${\mathscr{M}}_{0,4}^{c}\cong{\mathbb{P}}^{1}$ is a fine moduli space with positive Euler characteristic. Nevertheless, the corollary does hold in this case, and the proof is elementary.
Proposition 8.6.
Let $X={\mathscr{M}}_{0,n_{1}}\times\ldots\times{\mathscr{M}}_{0,n_{r}}$, where all $n_{i}\geq 3$. If $P\in\operatorname{Perv}(X)$, then $\chi(X,P)\geq 0$.
Proof.
Since ${\mathscr{M}}_{0,3}$ is a point, we may assume that $n_{i}>3$. We note that ${\mathscr{M}}_{0,4}\cong{\mathbb{P}}^{1}-\{0,1,\infty\}$. If $E=0+1+\infty$, then $\Omega_{{\mathscr{M}}_{0,4}}^{1}(\log E)$ is nef. Therefore the proposition holds for $X=({\mathscr{M}}_{0,4})^{r}$ by lemmas 4.4 and 4.5. We have an open immersion $j:{\mathscr{M}}_{0,n}\to({\mathscr{M}}_{0,4})^{n-3}$ sending
$(C,p_{1},\ldots,p_{n})\mapsto((C,p_{1},p_{2},p_{3},p_{i}))_{i=4,\ldots n}$
[…]
is also an open affine embedding. Therefore $J_{*}P$ is perverse when $P\in\operatorname{Perv}(X)$. Therefore $\chi(X,P)\geq 0$. ∎
Remark 8.7.
We note that Theorem 1.2 stated in the introduction follows from Theorem 5.3, Corollary 8.4 and Proposition 8.6.
@@
In this section, we discuss some pecularities in the characteristic $p$ setting. In particular, we show that the analog of Theorem 5.2 is false in this setting. We fix $N\geq 3$, and consider the moduli space $\prescript{}{N}{{\mathscr{A}}}_{g}$, viewed as a scheme over $\operatorname{Spec}{\mathbb{Z}}[1/N]$. Recall, that this is a smooth scheme over the base. Let $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}$ (resp. $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{C}}$) denote the corresponding moduli space over ${\mathbb{F}}_{p}$ (resp. ${\mathbb{C}}$). For a fixed prime $p$, let $\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p})$ denote the usual $\ell$-adic Euler characteristic (over $\bar{{\mathbb{F}}}_{p}$) for some fixed prime $\ell\neq p$. Note that this is and integer independent of $\ell$. Similarly, let $\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{C}})$ denote the usual topological Euler characteristic (say with rational coefficients). Note that this can be identified with the $\ell$-adic Euler charactersitic by the usual comparison theorem for étale cohomology and singular cohomology. We begin by first observing the following:
@@
Choose a good compactification (over ${\mathbb{Z}}[1/N]$) (for example, a toroidal compactification as in [10]) and use the lemma below. ∎
@@
Let $S$ be a strictly henselian ring with with closed point $s$ and generic point $\eta$. Let $\bar{X}\rightarrow S$ a smooth proper morphism, $D\subset\bar{X}$ a relative normal crossings divisor, and $X:=X\setminus D$. Then $\chi(X_{s})=\chi(X_{\bar{\eta}})$.
@@
Note that the analogous assertion for $\bar{X}$ holds as a corollary to the classical smooth and proper base change theorem. Suppose $D=D_{1}\cup\cdots\cup D_{n}$ is a decomposition into irreducible component, and let $D_{I}:=\bigcap_{i\in I}D_{i}$ for $I\subset\{1,\ldots,n\}$. Then each $D_{I}\rightarrow S$ is smooth proper, and therefore one also has $\chi((D_{I})_{s})=\chi((D_{I})_{\bar{\eta}})$. By descent, we conclude that $\chi(D_{s})=\chi(D_{\bar{\eta}})$. Here we use the stronger property that the Betti numbers are the same. Finally, by additivity of Euler characteristics we get the desired conclusion for $X$. ∎
@@
We note that for any smooth subvariety of $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}$ whose closure has normal crossings with $D$ (so that in particular the closure is smooth) and lifts to characteristic 0, we obtain the signed Euler char property as a consequence of the previous Lemma.
We now recall some counter-examples in characteristic p resulting from super-singular strata. Below, for a fixed prime $p$, we set $\prescript{}{N}{{\mathscr{A}}}_{g,p}:=\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}[\zeta_{N}]$ (where $\zeta_{N}$ is a fixed primitive $N$-th root of unity). Note that if $N=1$, we will denote this simply by ${\mathscr{A}}_{g,p}$. We denote by $|\prescript{}{N}{{\mathscr{A}}}_{g,p}|$ the corresponding coarse modulii spaces. If $N>2$, then this is the same as the fine moduli space.
Let $\prescript{}{N}{{\mathcal{S}}}_{g,p}\subset|\prescript{}{N}{{\mathscr{A}}}_{g,p}|$ denote the super singular locus. By definition this is the locus of points $(A,\lambda,H)$ (here $\lambda$ is a principal polarization and $H$ is a level structure) such that $A\otimes\bar{{\mathbb{F}}}_{p}$ is supersingular. This is a closed subset with $\prescript{}{N}{\dim}({\mathcal{S}}_{g,p})=\lfloor\frac{g^{2}}{4}\rfloor$ ([22] ). Note that it is known that in general $|{\mathcal{S}}_{g,p}|$ is not reducible, but it is equidimensional. In particular, $\dim(\prescript{}{N}{{\mathcal{S}}}_{2,p})=1$ and $\dim({\mathcal{S}}_{3,p})=2$. The components of ${\mathcal{S}}_{g,p}$ can be parametrized by spaces flag type data for Dieudonne modules ([22]). We do not recall the details, but give a description of the resulting parameter space in the case $g=2$ ([17] (section 2, 5.1, 5.3) and $g=3$ [22]
@@
    There is a smooth projective curve $C\subset\prescript{}{N}{{\mathscr{A}}}_{2,p}$ which does not satisfy the signed Euler characteristic property i.e. the $\ell$-adic Euler charactersitic $\chi(C_{\bar{{\mathbb{F}}}_{p}})>0$. In fact, $C={\mathbb{P}}^{1}$.
@@
The first part is a direct consequence of the first part of the previous theorem. For the second part, we note that the blow down of the surface given in part two satisfies the properties stated here. Namely, it is a normal surface with isolated singular point, and it has negative Euler characteristic. This follows from the fact that the Fermat curve in this case has negative Euler characteristic, and therefore so does the projective bundle described above. ∎
@@
    There is an embedding $({\mathbb{P}}^{1})^{[g/2]}\subset\prescript{}{N}{{\mathscr{A}}}_{g,p}$ for all $g\geq 2$.
@@
If $g=2$, (1) is the first part of the previous theorem. Suppose $g>2$, and let $d=[g/2]$. If $g$ is odd, let $E$ be a fixed polarized elliptic curve with level $N$ structure. Then an irreducible component of either
@@
e.g. by [10, chap IV, theorem 7.7]. Therefore $F^{1}$ cannot be nef.
@@
    The corollary gives a counterexample to Theorem 2.2 on the semipositivity (i.e. nefness) of Hodge bundles in characteristic p. This goes back to Moret-Bailly [25] who used essentially the same example when $g=2$.

### v2

@@ 7. Euler characteristics of perverse sheaves on ${\mathscr{M}}_{g,n}$
For a product ${\mathscr{M}}=\prod{\mathscr{M}}_{g_{i},n_{i}}^{sc}$ as in the theorem, let $H({\mathscr{M}})=\max_{i}d_{i}$ and $S({\mathscr{M}})=\Sigma_{i}d_{i}$ where $d_{i}=3g_{i}-3+n_{i}$ (the dimension of the given component). We consider pairs of natural numbers $(H,S)$ ordered lexicographically: $(H,S)<(H^{\prime},S^{\prime})$ if $H<H^{\prime}$ or $H=H^{\prime}$ and $S<S^{\prime}$. We shall proceed by induction with respect to this ordering.
Fix $N\geq 3$, and let ${\mathscr{M}}={}_{N}{\mathscr{M}}_{g_{1},n_{1}}^{sc}\times\ldots\times{}_{N}{\mathscr{M}}_{g_{r},n_{r}}^{sc}$. Here, we set $_{N}{\mathscr{M}}_{0,n}^{sc}:={}_{N}{\mathscr{M}}_{0,n}$. It is enough to prove that
(7.5.1) $\chi({\mathscr{M}},K)\geq 0$
for $K\in\operatorname{Perv}({\mathscr{M}})$.
We set $_{N}{\mathscr{A}}_{0,n}:={}_{N}{\mathscr{M}}_{0,n}$. Then, we have the induced map $\tau:{\mathscr{M}}\rightarrow{\mathscr{A}}$ where ${\mathscr{A}}=\prod{}_{N}{\mathscr{A}}_{g_{i},n_{i}}$ and $\tau$ given by the product of the identities on $_{N}{\mathscr{M}}_{0,n}$ and extended Torelli maps on other factors. Note that $\mathscr{A}$} has $PEC$ by Theorem 4.2. Since ${\mathscr{Y}}=\tau({\mathscr{M}})\subset{\mathscr{A}}$ is closed by Lemma 6.4, it also has PEC. Let $\mathscr{Z}$$\subset{\mathscr{M}}$ be the closed subset defined as the union of the closed subsets of $\mathscr{M}$} given by replacing the $i$-th factor in $\mathscr{M}$} by a divisor in the boundary ${\mathscr{M}}_{g_{i},n_{i}}^{sc}-{\mathscr{M}}_{g_{i},n_{i}}$. Then $\mathscr{Z}$} empty if and only if all the factors are either ${\mathscr{M}}_{0,n}$ or ${\mathscr{M}}_{1,1}$. In this case, we have already observed the conclusion of the theorem in the previous proposition. Therefore we can assume that $\mathscr{Z}$} is a nonempty divisor. By an application of Theorem 5.1, we are reduced to showing the $PEC$ property on the boundary $\mathscr{Z}$}.
The boundary $\mathscr{Z}$} is a union of divisors $\mathscr{D}$}, each of which is of the above form (a product of spaces $_{N}{\mathscr{M}}_{g^{\prime},n^{\prime}}^{sc}$). For each component $\mathscr{D}$}, we claim that $(H({\mathscr{D}}),S({\mathscr{D}}))<(H({\mathscr{M}}),S({\mathscr{M}}))$. To see this, note that a boundary divisor $\mathscr{D}$} amounts to replacing one of the factors in $\mathscr{M}$} by one of boundary divisors of the given factor. In particular, $S$ for this factor is one less (it is the dimension of this divisor). On the other hand, $H$ is the maximum dimension of the factors, and this either decreases or remains the same. This proves the claim.
It remains to show that $\chi({\mathscr{Z}},K)\geq 0$ for $K\in\operatorname{Perv}({\mathscr{Z}})$. On the other hand, since $\operatorname{Perv}({\mathscr{Z}})$ is Artinian, we can assume that $K$ is simple. Then $K$ is the image of a perverse sheaf $L$ in some ${\mathscr{Z}}^{\prime}=\prod{}_{N}{\mathscr{M}}_{g_{i}^{\prime},n_{i}^{\prime}}$, with $(H({\mathscr{Z}}),S({\mathscr{Z}}))<(H({\mathscr{M}}),S({\mathscr{M}}))$. Therefore, the claim follows by the induction hypothesis.
As an immediate consequence we have the following corollary, which completes the proof of Theorem 1.2.
@@
Let ${\mathscr{M}}^{sc}={\mathscr{M}}_{g_{1},n_{1}}^{sc}\times\ldots\times{\mathscr{M}}_{g_{r},n_{r}}^{sc}$. The embedding $j:{\mathscr{M}}\to{\mathscr{M}}^{sc}$ is affine, since the complement is a divisor. Therefore $j_{*}P\in\operatorname{Perv}({\mathscr{M}}^{sc})$. Consequently,
$\chi^{orb}({\mathscr{M}},P)=\chi^{orb}({\mathscr{M}}^{sc},j_{*}P)\geq 0$
@@
In this section, we discuss some peculiarities in the characteristic $p$ setting. In particular, we show that the analog of Theorem 4.2 is false in this setting. We fix $N\geq 3$, and consider the moduli space $\prescript{}{N}{{\mathscr{A}}}_{g}$, viewed as a scheme over $\operatorname{Spec}{\mathbb{Z}}[1/N]$. Recall, that this is a smooth scheme over the base. Let $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}$ (resp. $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{C}}$) denote the corresponding moduli space over ${\mathbb{F}}_{p}$ (resp. $\mathbb{C}$}). For a fixed prime $p$, let $\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p})$ denote the usual $\ell$-adic Euler characteristic (over $\bar{{\mathbb{F}}}_{p}$) for some fixed prime $\ell\neq p$. Note that this is an integer independent of $\ell$. Similarly, let $\chi(\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{C}})$ denote the topological Euler characteristic (say with rational coefficients). Note that this can be identified with the $\ell$-adic Euler characteristic by the usual comparison theorem for étale cohomology and singular cohomology. We begin by first observing the following:
@@
Choose a good compactification (over ${\mathbb{Z}}[1/N]$) (for example, a toroidal compactification as in [fc]) and use the lemma below. ∎
@@
Let $S$ be a strictly henselian ring with with closed point $s$ and generic point $\eta$. Let $\bar{X}\rightarrow S$ a smooth proper morphism, $D\subset\bar{X}$ a relative normal crossings divisor, and $X:=\bar{X}\setminus D$. Then $\chi(X_{s})=\chi(X_{\bar{\eta}})$.
@@
Note that the analogous assertion for $\bar{X}$ holds as a corollary to the classical smooth and proper base change theorem. Suppose $D=D_{1}\cup\cdots\cup D_{n}$ is a decomposition into irreducible component, and let $D_{I}:=\bigcap_{i\in I}D_{i}$ for $I\subset\{1,\ldots,n\}$. Then each $D_{I}\rightarrow S$ is smooth proper, and therefore one also has $\chi((D_{I})_{s})=\chi((D_{I})_{\bar{\eta}})$. By desecent, we conclude that $\chi(D_{s})=\chi(D_{\bar{\eta}})$. Alternatively, we can also conclude via the inclusion-exclusion principle. Finally, by additivity of Euler characteristics we get the desired conclusion for $X$. ∎
@@
Consider a smooth compactification of $\prescript{}{N}{{\mathscr{A}}}_{g}$ with normal crossings boundary $\mathscr{D}$} inducing a smooth compactification of $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}$ with boundary $D$. For any smooth subvariety of $\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}$ whose closure has normal crossings with $D$ (so that in particular the closure is smooth) and lifts to characteristic 0, we obtain the signed Euler characteristic property as a consequence of the previous Lemma.
We now recall some counter-examples in characteristic p resulting from super-singular strata. For the rest of this section, we assume that $p$ does not divide $N$. Below, we set $\prescript{}{N}{{\mathscr{A}}}_{g,p}:=\prescript{}{N}{{\mathscr{A}}}_{g}\otimes{\mathbb{F}}_{p}[\zeta_{N}]$ (where $\zeta_{N}$ is a fixed primitive $N$-th root of unity). Below, if $N=1$, we will sometimes denote $\prescript{}{N}{{\mathscr{A}}}_{g,p}$ simply by ${\mathscr{A}}_{g,p}$. We denote by $|\prescript{}{N}{{\mathscr{A}}}_{g,p}|$ the corresponding coarse modulii spaces. If $N>2$, then this is the same as the fine moduli space.
Let $\prescript{}{N}{{\mathcal{S}}}_{g,p}\subset|\prescript{}{N}{{\mathscr{A}}}_{g,p}|$ denote the super-singular locus. If $N=1$, we drop the superscript and denote it by ${\mathcal{S}}_{g,p}$. By definition this is the locus of points $(A,\lambda,H)$ (here $\lambda$ is a principal polarization and $H$ is a level structure) such that $A\otimes\bar{{\mathbb{F}}}_{p}$ is super-singular. This is a closed subset with $\dim(\prescript{}{N}{{\mathcal{S}}}_{g,p})=\lfloor\frac{g^{2}}{4}\rfloor$ ([lioort] ). Note that it is known that in general $|\prescript{}{N}{{\mathcal{S}}}_{g,p}|$ is not irreducible, but it is equidimensional. In particular, $\dim(\prescript{}{N}{{\mathcal{S}}}_{2,p})=1$ and $\dim(\prescript{}{N}{{\mathcal{S}}}_{3,p})=2$. The components of $\prescript{}{N}{{\mathcal{S}}}_{g,p}$ can be parametrized by spaces of flag type data for Dieudonne modules ([lioort]). We do not recall the details, but give a description of the resulting parameter space in the case $g=2$ ([KO] (section 2, 5.1, 5.3) and $g=3$ [lioort]
@@
    There is a smooth projective curve $C\subset\prescript{}{N}{{\mathscr{A}}}_{2,p}$ which does not satisfy the signed Euler characteristic property i.e. the $\ell$-adic Euler characteristic $\chi(C_{\bar{{\mathbb{F}}}_{p}})>0$. In fact, $C={\mathbb{P}}^{1}$.
@@
The first part is a direct consequence of the first part of the previous theorem. For the second part, we note that the blow down of the surface $\mathcal{Y}$} (i.e. $\pi({\mathcal{Y}})$) in (2) of Theorem 8.5 satisfies the properties stated here. Namely, it is a normal surface with isolated singular point, and it has negative Euler characteristic. This follows from the fact that the Fermat curve in this case has negative Euler characteristic, and therefore so does the projective bundle described above. ∎
@@
    There is an embedding $({\mathbb{P}}^{1})^{\lfloor g/2\rfloor}\subset\prescript{}{N}{{\mathscr{A}}}_{g,p}$ for all $g\geq 2$.
@@
If $g=2$, (1) is the first part of the previous theorem. Suppose $g>2$, and let $d=\lfloor g/2\rfloor$. If $g$ is odd, let $E$ be a fixed polarized elliptic curve with level $N$ structure. Then an irreducible component of either
@@
e.g. by [fc, chap IV, theorem 7.7]. Therefore $F^{1}$ cannot be nef.
@@
    The corollary gives a counterexample to Theorem 2.4 on the semipositivity (i.e. nefness) of Hodge bundles in characteristic p. This goes back to Moret-Bailly [mb] who used essentially the same example when $g=2$.